OpenStudy (anonymous):

A French class has a total of 42 students. The number of males is 6 more than the number of females. How many males and how many females are in the class?

sammixboo (sammixboo):

Whoops hold on wrote it wrong

sammixboo (sammixboo):

x = males x + (x - 6) = 42 Right @sleepyjess

sammixboo (sammixboo):

I am confusing myself, haha

sammixboo (sammixboo):

Or would it be x = females x + (x + 6) = 42

sammixboo (sammixboo):

I mean they are the same I guess

OpenStudy (sleepyjess):

f = x m = x + 6 Combine, x + x + 6 = 42

sammixboo (sammixboo):

Ok. :3 So do you know what x is @Deldrickg

OpenStudy (sleepyjess):

The first was Sammi did it would also be fine. m = x f = x - 6 x + x - 6 Either way you choose is fine :)

OpenStudy (anonymous):

is x 7

OpenStudy (sleepyjess):

For which set of equations?

OpenStudy (sleepyjess):

Either way, it wouldn't work out.

OpenStudy (sleepyjess):

But we need to know which equation you want to use.

sammixboo (sammixboo):

Let's just stick with x + (x + 6) = 42

OpenStudy (anonymous):

Ok,

sammixboo (sammixboo):

So to start solving this, you combine the two x's to get 2x + 6 = 42. Now can you give another shoot at solving for x?

OpenStudy (anonymous):

yes 2x+6=42 -2x -2x

sammixboo (sammixboo):

Not quite. You wanna start with the 6 not the 2x

OpenStudy (sleepyjess):

Close, we need to keep x on the left and move + 6 to the right

OpenStudy (sleepyjess):

*keep 2x

sammixboo (sammixboo):

So subtract 6 instead of 2x

sammixboo (sammixboo):

2x + 6 - 6 = 42 - 6 2x = 42 - 6

OpenStudy (anonymous):

2x=36

sammixboo (sammixboo):

Right! Now do you know what x is?

OpenStudy (anonymous):

x=18

sammixboo (sammixboo):

Right :)

sammixboo (sammixboo):

18 + (18 + 6) = 42 18 + 24 = 42 42 = 42

OpenStudy (anonymous):

And this is for Males and Females

sammixboo (sammixboo):

Oh since x = females, there is 18 females

sammixboo (sammixboo):

And since there is 6 more males than females, how many males would that be?

OpenStudy (anonymous):

ok24

sammixboo (sammixboo):

f = x m = x + 6

sammixboo (sammixboo):

Right!